OpenStudy (anonymous):

solve the inequality. - l 2x-5 l greater than or equal to -9 i got (-2,7) but it says its wrong

OpenStudy (anonymous):

ok first thing we need to do is isolate the absolute value so we would divide -9 by - which will turn it to positive 9

OpenStudy (anonymous):

2x-5 less than or equal to 9 2x-5 greater than or equal to -9 can you solve this

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

2x less than or equal to 4

OpenStudy (anonymous):

positive or negative?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

which one are you doing right now the 2nd one?

OpenStudy (anonymous):

well its both

OpenStudy (anonymous):

\[2x \le 14\] \[2x \ge -4\] are the equations i got before isolating x

OpenStudy (anonymous):

oo ok

OpenStudy (anonymous):

now can you isolate x

OpenStudy (anonymous):

x less than or equal to 7 and x is less than or equal to -2

OpenStudy (anonymous):

\[x \le 7\] \[x \ge -2\] now can you combine the ranges

OpenStudy (anonymous):

-2≤x≤7

OpenStudy (anonymous):

\[-2 \le x \le 7 \]

OpenStudy (anonymous):

-2 less than or equal to x is less than or equal to 7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can i write (-2,7)?

OpenStudy (anonymous):

you cant use this one?

OpenStudy (anonymous):

no it won't let me

OpenStudy (anonymous):

hmm im not sure bud they might be wrong because i just used a calculator for this and it got the same thing id ask your teacher

OpenStudy (anonymous):

thank you!